Seven years ago, Yves was six times as old as Eric. in one year, he will only be twice as old as Eric. What are their present ages?: Seven years ago, Yves was six times as old as Eric. in one year, he will only be twice as old as Eric. What are their present ages?
Answer by stanbon(17981)

Seven years ago, Yves was six times as old as Eric. in one year, he will only be twice as old as Eric. What are their present ages?
---------------------
Let Yves be y yrs. old now.
Let Eric be e yrs. old now.
------------------------------
EQUATIONS:
(y-7) = 6(e-7)
y+1 = 2(e+1)
------------------
Rearrange equations:
y = 6e -35
y = 2e + 1
---------------
6e -35 = 2e+1
4e = 36
e = 9 yrs old (Eric's age now)
---------
y = 2*9+1 = 19 (Yves's age now)
=====================
Cheers,
Stan H.